17. Graph the system of equation shown below. Then solve the system to find the points of intersection. {2+ y2 + 4y = 0
Answer #1

Date Page syste Given, tem o equation 212+y2 = 5 T x² + y² +4y=0 Now, 22.4.42 +44 = 9 =) 8t4y=8 (* *?442 = 8 ) lie, y-2 i,ep